OpenStudy (anonymous):

integrate sin^2x / sqrt x dx

OpenStudy (anonymous):

We'll replace (sin x)^2 = (1-cos 2x)/2 Int (1-cos 2x)dx/2sqrt x = Int dx/2sqrt x - Int cos 2x dx/2sqrt x Int (1-cos 2x)dx/2sqrt x = sqrt x - Int cos 2x dx/2sqrt x We'll integrate by parts Int cos 2x dx/2sqrt x using formula Int udv = u*v - Int vdu

OpenStudy (dumbcow):

that still won't work http://www.numberempire.com/integralcalculator.php?function=cos%282x%29%2F%282*sqrt%28x%29%29&var=x&answers=&__utma=164170016.165112442.1306893325.1311847678.1311849610.5&__utmz=164170016.1311849611.5.4.utmcsr%3Dgoogle |utmccn%3D%28organic%29|utmcmd%3Dorganic|utmctr%3Dintegral+calculator&screen_height=768&__utmc=164170016&__utmb=164170016.3.10.1311849611

OpenStudy (anonymous):

This seems to be one of those u keep getting an integral that is as difficult as the one u started with. Let me look around, see if there is some tricky way around it.

OpenStudy (anonymous):

ok :(

OpenStudy (anonymous):

i am having a test and this is a hint

OpenStudy (anonymous):

Normally u would do integration by parts twice and then use the result of the second integration to subback which I am just having a look at now but seems I have to go at least 3 times....:-(

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I can't be doing all the calculations (doing my head in) but if u use georgia's substitution and integrate by parts twice u will go from Cos 2x to Sin 2x and back to Cos 2x so u should be able to substitute from the first to the last for an answer (I think) Else u will have to wait for the calc wiz's to arrive, sorry...

OpenStudy (dumbcow):

im telling you there is no solution using elementary functions, the anti-derivative involves the error-function

OpenStudy (anonymous):

Strange they would give such a thing in a test, though...

OpenStudy (anonymous):

i hope not... he said::: could be in the test

OpenStudy (anonymous):

Multiple choice? D-no solution...:-)

OpenStudy (anonymous):

jajajaja

OpenStudy (dumbcow):

hmm i dunno...you have a point though is there more to this problem?

OpenStudy (anonymous):

what u mean?

OpenStudy (anonymous):

I am going to have to try what I said, I hate number crunching. Where's satellite, he loves this sort of stuff...

OpenStudy (dumbcow):

is it a definite integral?

OpenStudy (anonymous):

i read to solve this problem they use rengels,, something like that

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

from 0 to pi

OpenStudy (dumbcow):

oh ok, then you can evaluate it using numerical methods. Even graph it in a graphing calculator then estimate area from 0 to pi

OpenStudy (anonymous):

Like 1 and a bit...

OpenStudy (dumbcow):

yeah...i get 1.34

OpenStudy (anonymous):

dumbcow looks right, the problem is the powers of x go up so there is no exact substitution that I can see. I don't know why they bother with all this stuff anyway, real world integrals never work like the ones in the books...:-) They are nearly always have to be solved numerically. Must be for pure math people.

OpenStudy (anonymous):

:(